# Gaia Data Release 3: spectroscopic binary-star orbital solutions

### v1

@@ Abstract
Context. The Gaia satellite constitutes one of the cornerstone missions of ESA. This is an astrometric space experiment that is measuring positions, proper motions as well as parallaxes for a huge number of stars. It is also performing photometric and spectrophotometric observations. Gaia also operates a medium-dispersion spectrometer, the RVS, that provides spectra and thus radial velocity (RV) time-series.
Aims. The paper is centred on the analysis of the RV time-series. We try to fit orbital models and trend models and restrict ourselves to the objects of spectral types F-G-K brighter than magnitude 12 presenting only one single spectrum (SB1).
Methods. Suitable time-series are processed and analysed object per object, providing orbital or trend solutions. The results of the various fits are further filtered internally on the basis of several quality measures in order to discard spurious solutions. The objects only having a spectroscopic solution are classified in one of the three classes SB1 (eccentric model), SB1C (circular model) or TrendSB1 (mere trend model).
Results. We detail the methods used and describe the derived parameters and results. After a description of the models considered, and of the related quality tests of the fit, we detail the internal filtering that is intended to reject bad solutions. We also present a full validation of the pipeline. A description of the current content of the catalogue is provided.
Conclusions. We present the SB1, SB1C and the TrendSB1 spectroscopic solutions contained in the SB-subcatalogue which is part of the DR3 catalogue. We delivered some 181327 orbital solutions in the class SB1, 202 in the class SB1C and 56808 in the associated class TrendSB1. This is a first release and the delivered SB-subcatalogue could be tuned and refined. However the majority of the entries are correct; this data set constitutes by far the largest set of spectroscopic orbital solutions ever constituted.
@@
The Gaia satellite constitutes one of the cornerstone missions of the European Space Agency. This is an astrometric space experiment intended to represent a new breakthrough compared to the previous pioneering one named Hipparcos (Perryman 2009).
@@
The Gaia satellite has two telescopes, looking at 1065 of each other, and three associated scientific instruments, continuously scanning the celestial sphere thanks to its six-hour spinning on itself (around an axis perpendicular to the two telescope lines of sight), to the precession of the spinning axis around the direction of a fictitious, nominal Sun (compensating for the orbit of the satellite around the Sun-Earth Lagrangian point L2) with a 63-day period and to the one-year orbit of the satellite around the Sun accompanying the Earth in its revolution (Gaia Collaboration et al. 2016b). Multiple visits are necessary to reach a convenient S/N ratio, but also to gather information on binarity (or higher order multiplicity) and variability.
@@
The Gaia satellite is measuring the positions and transverse proper motions for some 1.8 billion stars, as well as precise parallaxes and thus distances for a very large number of objects. The detectors are working in a Time Delay Integration mode (TDI). The satellite was launched on December 19th, 2013 and is scanning the sky since the end of the commissioning period in July 2014. The detailed description of the project and of the payload can be found in Gaia Collaboration et al. (2016b) that was published along with the first Data Release (DR1). The description of DR1 and of the related data products is provided in Gaia Collaboration et al. (2016a) whereas the DR2 was detailed in Gaia Collaboration et al. (2018a). The present paper is to be associated with the DR3 (Gaia Collaboration et al. 2023d). Besides astrometry performed with the Astrometric Field (AF) instrument (Lindegren et al. 2018, 2021), the satellite also produces photometry from the same instrument as well as spectrophotometry provided by another channel: the BP/RP spectrophotometer purveyor of very low resolution spectra as well as of photometric colours (Riello et al. 2018, 2021; Evans et al. 2018). The two above-mentioned channels are supposed to observe stellar objects down to about magnitude 20th in band.
@@
Besides astrometry and spectrophotometry, the satellite also incorporates a third channel: a medium resolution spectrometer to measure the radial velocities (RVs) of stellar objects, named the Radial Velocity Spectrometer (RVS). This channel and the related instrument are described in great details in Cropper et al. (2018) which also includes a nice review of the genesis of the RVS instrument. The RVS brings the third component of the space velocity (to join with the transverse proper motion on the sky). It also provides for the brightest stars information on the nature of the object (, , chemical abundances, chromospheric activity, and so on). Stars brighter than magnitude = 2.76 produce saturated pixels and are not considered (for the definition of the magnitudes, see Sartoretti et al. 2023).
@@
The scanning law for RVS is intended to typically produce 40 transits over the nominal five-year mission. The focal plane of the RVS is covered by 12 CCDs (4 rows of 3 in a row). The RVS instrument acquires, in TDI mode, spectra in the range 845-872 nm, including the well-known Ca ii triplet (in late-type stars). With a resolving power of about (see section 8.2 and 11 of Cropper et al. 2018), it obtains three spectra per transit, each corresponding to a maximum exposure time of 4.42 s (they are analysed individually and concomitantly). These spectra are spread over about 1100 pixels (Cropper et al. 2018), but the part used for RV determination is restricted to the range 846-870 nm (Sartoretti et al. 2023). Very bright objects benefit from 2D spectra whose across scan component is 10 pixels.
@@
The data coming out of the RVS are treated by ground-based spectroscopic processing. The spectroscopic processing is ensured by the Coordination Unit 6 (CU6) of the Data Processing and Analysis Consortium (DPAC; Gaia Collaboration et al. 2016b). The first description of this processing was given in Sartoretti et al. (2018) for DR2 and the properties and validation of the RV set were described in Katz et al. (2019). The stars exhibiting a double-line spectrum (also named composite spectrum) were rejected from the process (as described in Sartoretti et al. 2018; Damerdji et al. 2024). The spectroscopic processing pipeline is made of a series of modules, each taking into account a series of tasks. The first modules are working per transit and per trending epochs. The flowchart is illustrated in section 2.3 of Sartoretti et al. (2018). The Single Transit Analysis (STA) is the workflow that actually performs the measurements of the RV per transit. Three submodules are measuring the RVs of all the selected convenient stars (single spectrum ones). They are based on the idea to derive the RV by cross-correlating the observed spectrum with a synthetic spectrum built in generate_template (Sartoretti et al. 2018). The interest to make use of several methods is explained in the early investigation of the problem, whose conclusions are detailed in David et al. (2014) (see also section 7 of Sartoretti et al. 2018). The STA measures the RVs per object and per transit. Thereafter, the Multiple Transit Analysis (MTA) workflow performs some statistical analysis tests based on the RVs from all objects at all transits. It computes per star a probability of constancy that is used to spot out variable objects. It also computes the median of the RVs over all transits for all the bright objects (Sartoretti et al. 2018). The constancy probability is not used to define the list of assumed constant objects that are considered to compute their median RV. The updated version of the spectroscopic chain for DR3 is described in Sartoretti et al. (2022); Katz et al. (2023). In DR3, MTA further computes mean spectra for assumed constant objects (Seabroke et al., in preparation).
@@
In the context of DR3, the RV values corresponding to variable objects are transmitted to the CU4 (object processing) and in particular to the Non-Single Star (NSS) part of the object processing that analyses these data in order to derive possible orbital solutions. Three channels running in parallel deal with the astrometric data (Halbwachs et al. 2023; Holl et al. 2023b), the photometric data (Siopis et al. 2024; Pourbaix et al. 2022, see section 7.6 by Siopis & Sadowski) and the spectroscopic data (the present paper, and Damerdji et al. 2024). The detailed description of the CU4 spectroscopic processing chain for SB1 (spectroscopic binaries with only one spectrum visible; hereafter labelled NSS-SB1) orbital solutions or trend ones is the subject of the present paper whereas Damerdji et al. (2024) describes the SB2 channel. These DR3 products are preliminary results; this is the first time that orbital solutions issued from Gaia are delivered. The results of the three channels can be further processed by combining their independent outputs. This is the task of the combiner (Gaia Collaboration et al. 2023a; Pourbaix et al. 2022, see section 7.7 by Gavras & Arenou). The scientific conception of the SB1 chain described here was elaborated by Yassine Damerdji, Eric Gosset and Thierry Morel; most of the codes were benchmarked in MATLAB and finally written in Java by YD with some help from TM; the validation tests (including the critical building of the ground-based data catalogues used for comparison) were mainly performed by TM. The whole work was made in the highly favourable and boosting context of the NSS team.
@@
Section 2 describes the nature of the data that are entering the process, whereas Section 3 describes the simple mathematical formalism used and lists the models that are considered. Section 4 explicits the importance of the period search and the related difficulties. Sect. 5 briefly summarises the main trials we did to finally arrive at the method described in Sect. 6 that presents the chain as adopted for the DR3 (NSS-SB1 main processing) whereas Sect. 7 introduces the notion of post-processing. Section 8 illustrates the results. Section 9 concerns the validation of the results of the chain. Additional considerations in Sect. 10 explain the details of the post-filtering and provide additional information in the light of the conclusions from the previous Section. The catalogue description, the caveats and the precautions necessary in using the results of the NSS-SB1 processing are given in Section 11. Section 12 concludes the paper. The content of the present paper is more detailed and contains material that supplants the information given in the Gaia archives (see sections 7.4 and 7.5 in Pourbaix et al. 2022)111https://gea.esac.esa.int/archive/documentation/GDR3/ at the epoch of the release.
@@ 2 Description of the input data
The DR3 data correspond to the first 34 months of the nominal mission (25 July 2014 - 28 May 2017, almost 1038 d, 2.84 y). The instrumental spectra obtained at each RVS transit are converted to proper physical spectra (reduced) by the pipeline developed by the CU6 (spectroscopic processing), see Sartoretti et al. (2022); Katz et al. (2023). Per transit, one spectrum is acquired by each of the three CCDs in a row of the RVS (Cropper et al. 2018). The RVs are measured individually on the three spectra. Three methods are used to derive the radial velocities (Cropper et al. 2018), each of them providing three velocities (one per CCD) and a combined RV value (over the three CCDs). The three combined RVs issued from the three methods are transmitted to the Integrator that provides a single combined value (per transit). Only this single combined value is considered here but the other intermediate RVs are used as a sanity check to validate the final combined value (see Sects. 6.4.8.2 and 6.4.8.5 of Sartoretti et al. 2022). The RVs are measured by cross-correlating the observed reduced and calibrated spectra with a theoretical/synthetic spectrum computed on the basis of various stellar atmosphere models. These so-called templates are identified through a set of three atmospheric parameters (effective temperature, surface gravity, metallicity). The selection of the template used to measure the RVs is based on these three parameters only; the alpha element content is not considered as an independent parameter. The assignation of the set of parameters to the object treated is performed on the basis of three databases. They are ordered here in the order of preference. The first possible origin is an internal compilation of ground-based catalogues. The second one is from an early run of the Gaia General Stellar Parametrizer from Spectroscopy (GSP-Spec, Recio-Blanco et al. 2023). The third one is from an early run of the Gaia General Stellar Parametrizer from Photometry (GSP-Phot, Andrae et al. 2023). Finally, if none of these sources provides the necessary parameters, the latter are estimated as previously done for DR2 (DetermineAP: section 6.5 of Sartoretti et al. 2018). This process DetermineAP cross-correlates the observed spectrum with each template from a limited subset of 28 selected ones, and this for every transit. The template giving per transit the highest correlation is retained, whatever is the velocity. The final adopted template is the one that is the major contributor over all the transits. More details can be found in section 3.6.1 of Katz et al. (2023). The repartition among the various origins for the data set treated here is the following: Ground-based catalogues (14.4 %), GSP-Spec (33.4 %), GSP-Phot (45.6 %) and DetermineAP (6.6 %). The RV measurements from transit to transit are performed with the same template. Consequently, a template mismatch is expected to essentially introduce a systematic RV shift (for single-line objects).
@@
The measurements of the RVs are performed under the responsibility of the STA Development Unit. The list of good RVs corresponding to a particular star are then analysed by the MTA in order to detect and separate variable RV sets from constant RV sets. For each star, the median of the RVs is computed and appears in the main catalogue. The stars with variable RVs are further considered by the global pipeline and are forwarded to the CU4 (Object processing: NSS). For DR3, the adopted cutting threshold corresponds to a probability of variability of 0.99 (rv_chisq_pvalue . The composite spectra and SB2 objects are separately treated already at the level of the CU6 processing (Damerdji et al. 2024) for the determination of both RVs and are then dispatched to CU4/NSS in order to search for the corresponding orbital solutions (Damerdji et al. 2024).
@@
The Non-Single Star (NSS) processing is intended to perform the reduction of the corresponding data from the point of view of multiple star studies. In the present paper, we will consider the sole spectroscopic channel for the SB1 objects.
@@
The input data are made of a list of RVs (and their 1-uncertainties) that are corrected to make reference to the barycentre of the solar system. It should be pointed out that for DR3 the rms error on the barycentric correction ( km s-1) is not propagated to the uncertainty on the individual RVs. The RVs are accompanied by a time of observations expressed in barycentric Julian days (BJD). The number of points of the time-series corresponds to the number of transits for which the RV has been properly measured with success. The data could present several weaknesses such that outliers and wrong RVs due e.g. to template mismatch errors (for the latter case, most probably the same systematic for all values within a time-series). These problems and additional selection rules are further documented in Katz et al. (2023) and will not be further detailed here. It should however be noticed that some objects have been rejected at the level of the input data filtering in CU6 because they correspond to a poorly performing Astrometric Global Iterative Solution (AGIS) solution or to a large error on the along-scan field angle (see its definition in Sartoretti et al. 2018). The threshold has been chosen at mas, equivalent to about 29 km s-1. Other filtering applied to the input CU6 data is described in section 4 of Katz et al. (2023). In the course of the CU6 STA/MTA validation, a bias on output RVs was noticed for stars fainter than (being 0 km s-1 at magnitude 11 and reaching 75 m s-1 at magnitude 12, Katz et al. 2023). The correction has not been applied for the catalogue, nor for the transmission of the RVs to the NSS processing. The effect is quite small and should only induce per object a global shift of all the RVs with as a consequence a small bias on the systemic velocities (see Sect. 3).
@@
The current astrometric solution for DR3 has been released as part of the EDR3 and contains, among others, positions for a large number of objects. The astrometric solution (AGIS) is based on a treatment assuming the objects are single, unresolved stars and thus having an image characterised by the instrumental line-spread function (for the one dimensional observations in the astrometric channel) or point-spread function (for the two-dimensional observations), see the description in Lindegren et al. (2021). Some objects do not behave as expected according to this assumption. The pathological behaviour could be due to anomalies, but a large contribution to this problem originates in unresolved or poorly resolved double (or multiple) objects. The separation threshold is considered to be around 04-05 (see Fabricius et al. 2021; Holl et al. 2023a). The double objects not following the behaviour for single stars are associated with a bad adjustment to the global astrometric solution and are thus accompanied by a large renormalised unit weighted error (ruwe). Lindegren et al. (2021) also associate to these objects an excess noise (impacting the position) that is delivered in the catalogue under the name astrometric_excess_noise and which corresponds to the additional dispersion necessary to accept the concerned object in the global solution. Objects with a ruwe larger than 1.4 are certainly problematic, but smaller values already have an impact on the RVS calibration. Since the RVS is a slitless spectrometer, the calibration in wavelength of the spectra is requesting an epoch position of the object that is accurate enough. Therefore, for the objects that are deviating from the single star behaviour, we could expect that the wavelength calibration is not correct, and thus the measured transit RVs are somewhat biased and this creates artefacts in the time-series. No correction of this effect is offered for DR3. As a first approach, we preferred not to discard these objects on the basis of the ruwe. In order to take into account this problem, it was decided to inflate the uncertainty on the RV value by introducing (quadratically) an additional component linked to the astrometry. This term is derived as being (km s-1) = 0.146 astrometric_excess_noise (mas) where the factor 0.146 is related to the dispersion of the spectrometer (transformation of mas into km s-1). Typically, the values of astrometric_excess_noise are below 10 mas, and thus is below 1.46 km s-1.
@@
Some selections are also made on the objects at the moment of their entry in the NSS-SB1 processing. In the case of DR3, the kind of stars is restricted to the range of effective temperatures 3875 K to 8125 K (mainly ((M))-K-G-F-(A) stars). No emission-line object has been selected, and in principle only stars without spectral peculiarities are entering the game. The objects with a rv_renormalised_gof larger than 4 were selected (Katz et al. 2023) to enter the NSS-SB1 processing because they were considered as clearly non-constant. The measurements of the individual (per transit) RVs are restricted to the range –1000 to +1000 km s-1. A global view of the selections applied beforehand is available in Table 18 (Appendix H). This measurement of the individual RVs on individual spectra is only possible for bright objects. The distribution in magnitude of the treated objects is provided in Fig. 15 in the validation Sect. 9.2.1. A threshold on has been selected at 12 mag. Fainter objects are considered to produce individual spectra for which the S/N ratio forbids to derive a sound radial velocity. This strongly reduces the amount of selected objects. The stellar RVs are not processed by the NSS spectroscopic channel if the number of transits (data points) is less than 10 (see details in Sect. 6.3.1).
@@
When the SB2 chain (Damerdji et al. 2024) analyses an SB2 time-series and is not able to derive the corresponding SB2 orbital solution, the object can be redirected to the SB1 chain. An SB1 time-series is built with the RVs characterised by the smallest uncertainties. The object is then considered as an SB1 and is forwarded to the present chain. This is a second possibility of entering the SB1 channel. These objects are flagged (for more details, see Damerdji et al. 2024, and also Appendix I).

### v2

@@ Abstract
Context. The Gaia satellite constitutes one of ESA’s cornerstone missions. Being primarily an astrometric space experiment measuring positions, proper motions, and parallaxes for a huge number of stars, it also performs photometric and spectrophotometric observations. Gaia operates a medium-dispersion spectrometer, known as Radial Velocity Spectrometer (RVS), which provides spectra and radial velocity (RV) time series.
Aims. The paper is focussed on the analysis of the RV time series. We fit orbital and trend models, restricting our study to objects of spectral types F-G-K that are brighter than a magnitude of 12, presenting only one single spectrum (SB1).
Methods. Suitable time series were processed and analysed on an object-per-object basis, providing orbital or trend solutions. The results of the various fits were further filtered internally on the basis of several quality measures to discard spurious solutions. The objects with solely a spectroscopic solution were classified in one of the three classes: SB1 (eccentric model), SB1C (circular model), or TrendSB1 (mere trend model).
Results. We detail the methods used in this work and describe the derived parameters and results. After a description of the models considered and the related quality tests of the fit, we detail the internal filtering process aimed at rejecting bad solutions. We also present a full validation of the pipeline. A description of the current content of the catalogue is also provided.
Conclusions. We present the SB1, SB1C, and TrendSB1 spectroscopic solutions contained in the SB subcatalogue, part of the DR3 catalogue. We deliver some 181 327 orbital solutions in class SB1, 202 in class SB1C, and 56 808 in the associated class TrendSB1. This is a first release and the delivered SB subcatalogue could be further tuned and refined. However, the majority of the entries are correct. Thus, this data set constitutes by far the largest set of spectroscopic orbital solutions to be computed.
@@
The Gaia satellite constitutes one of the cornerstone missions of the European Space Agency (ESA). This astrometric space experiment is intended to represent a new breakthrough, compared to the previous pioneering one, Hipparcos (Perryman 2009). The Gaia satellite has two telescopes, observing at 1065 of each other, and three associated scientific instruments that are continuously scanning the celestial sphere thanks to its six-hour rotation (around an axis perpendicular to the two telescope’s lines of sight), as well as the precession of the spinning axis around the direction of a fictitious, nominal Sun (compensating for the orbit of the satellite around the Sun-Earth Lagrangian point L2), with a 63-day period, and thanks to the one-year orbit of the satellite around the Sun, while it accompanies the Earth in its revolution (Gaia Collaboration et al. 2016b). Multiple visits are necessary to reach a convenient signal-to-noise (S/N) ratio, but also to gather information on binarity (or higher order multiplicity) and variability.
@@
The Gaia satellite is measuring the positions and transverse proper motions for some 1.8 billion stars, as well as precise parallaxes and thus distances for a very large number of objects. The detectors are working in a Time Delay Integration mode (TDI). The satellite was launched on 19 December 2013, and has been scanning the sky since the end of the commissioning period in July 2014. The detailed description of the project and of the payload can be found in Gaia Collaboration et al. (2016b), published alongside the first Data Release (DR1). The description of DR1 and of the related data products is provided in Gaia Collaboration et al. (2016a) whereas DR2 was detailed in Gaia Collaboration et al. (2018a). The present paper is to be associated with the DR3 (Gaia Collaboration et al. 2023d). In addition to astrometry performed with the Astrometric Field (AF) instrument (Lindegren et al. 2018, 2021), the satellite also produces photometry from the same instrument as well as spectrophotometry provided by another channel: the BP/RP spectrophotometer purveyor of very low-resolution spectra as well as of photometric colours (Riello et al. 2018, 2021; Evans et al. 2018). The two above-mentioned channels are supposed to observe stellar objects down to a magnitude of about 20 in the band.
@@
In addition to astrometry and spectrophotometry, the satellite also incorporates a third channel: a medium-resolution spectrometer to measure the radial velocities (RVs) of stellar objects, known as the Radial Velocity Spectrometer (RVS). This channel and the related instrument are described in great detail in Cropper et al. (2018), which also includes a comprehensive review of the genesis of the RVS instrument. The RVS brings the third component of the space velocity (to join with the transverse proper motion on the sky). It also provides, for the brightest stars, information on the nature of the object (, , chemical abundances, chromospheric activity, etc.). Stars brighter than a magnitude of = 2.76 produce saturated pixels and, thus, are not considered here (for the definition of the magnitudes, see Sartoretti et al. 2023).
@@
The scanning law for RVS is intended to typically produce 40 transits over the nominal five-year mission. The focal plane of the RVS is covered by 12 CCDs (four rows of three in a row). The RVS instrument acquires, in TDI mode, spectra in the range 845-872 nm, including the well-known Ca ii triplet (in late-type stars). With a resolving power of about (see Sects. 8.2 and 11 of Cropper et al. 2018), it obtains three spectra per transit, each corresponding to a maximum exposure time of 4.42 s (they are analysed individually and concomitantly). These spectra are spread over about 1100 pixels (Cropper et al. 2018), but the part used for RV determination is restricted to the range 846-870 nm (Sartoretti et al. 2023). Very bright objects benefit from 2D spectra whose across scan component is 10 pixels.
@@
The data coming out of the RVS are treated by ground-based spectroscopic processing. The spectroscopic processing is ensured by the Coordination Unit 6 (CU6) of the Data Processing and Analysis Consortium (DPAC; Gaia Collaboration et al. 2016b). The first description of this processing was given in Sartoretti et al. (2018) for DR2 and the properties and validation of the RV set were described in Katz et al. (2019). The stars exhibiting a double-line spectrum (also known as the composite spectrum) were rejected from the process (as described in Sartoretti et al. 2018; Damerdji et al. 2024). The spectroscopic processing pipeline is made of a series of modules, each taking into account a series of tasks. The first modules are working per transit and per trending epochs. The flowchart is illustrated in Sect. 2.3 of Sartoretti et al. (2018). The single transit analysis (STA) refers to the workflow that actually performs the measurements of the RV per transit. Three submodules are measuring the RVs of all the selected convenient stars (single spectrum ones). They are based on the idea to derive the RV by cross-correlating the observed spectrum with a synthetic spectrum built in generate_template (Sartoretti et al. 2018). The incentive in making use of several methods is explained in the early investigation of the problem, with conclusions detailed in David et al. (2014) (see also Sect. 7 of Sartoretti et al. 2018). The STA measures the RVs per object and per transit. Thereafter, the multiple transit analysis (MTA) workflow performs some statistical analysis tests based on the RVs from all objects at all transits. It computes (per star) a probability of constancy that is used to spot out variable objects. It also computes the median of the RVs over all transits for all the bright objects (Sartoretti et al. 2018). The constancy probability is not used to define the list of assumed constant objects that are considered to compute their median RV. The updated version of the spectroscopic chain for DR3 is described in Sartoretti et al. (2022); Katz et al. (2023). In DR3, MTA further computes mean spectra for assumed constant objects (Seabroke et al., in preparation).
@@
In the context of DR3, the RV values corresponding to variable objects are transmitted to the CU4 (object processing) and in particular to the non-single star (NSS) part of the object processing that analyses these data in order to derive possible orbital solutions. Three channels running in parallel deal with the astrometric data (Halbwachs et al. 2023; Holl et al. 2023b), photometric data (Siopis et al. 2024; Pourbaix et al. 2022, see Sect. 7.6 by Siopis & Sadowski), and the spectroscopic data (the present paper, and Damerdji et al. 2024). The detailed description of the CU4 spectroscopic processing chain for SB1 (spectroscopic binaries with only one spectrum visible; hereafter labelled NSS-SB1) orbital solutions or trend ones is the subject of the present paper whereas Damerdji et al. (2024) describe the SB2 channel. These DR3 products are preliminary results; this is the first time that orbital solutions issued from Gaia have been delivered. The results of the three channels can be further processed by combining their independent outputs. This is the task of the combiner (Gaia Collaboration et al. 2023a; Pourbaix et al. 2022, see Sect. 7.7 by Gavras & Arenou). The scientific conception of the SB1 chain described here was elaborated on by Yassine Damerdji, Eric Gosset and Thierry Morel; most of the codes were benchmarked in MATLAB and finally written in Java by YD with some help from TM; the validation tests (including the critical building of the ground-based data catalogues used for comparison) were mainly performed by TM. The whole work was carried out in the highly favourable and boosting context of the NSS team.
@@
Section 2 describes the nature of the data that are entering the process, whereas Sect. 3 describes the simple mathematical formalism used and lists the models that are considered. Section 4 explicits the importance of the period search and the related difficulties. Section 5 briefly summarises the main trials we did to arrive at the method described in Sect. 6, where we present the chain as adopted for the DR3 (NSS-SB1 main processing). In Sect. 7, we introduce the notion of post-processing. Section 8 illustrates the results. Section 9 concerns the validation of the results of the chain. Additional considerations in Sect. 10 explain the details of the post-filtering and provide additional information in the light of the conclusions from the previous section. The catalogue description, the caveats and the precautions necessary in using the results of the NSS-SB1 processing are given in Sect. 11. Section 12 presents our conclusions. The contents of the present paper are more detailed and contain material that supplants the information given in the Gaia archives (see Sects. 7.4 and 7.5 in Pourbaix et al. 2022)111https://gea.esac.esa.int/archive/documentation/GDR3/ at the time of the release. The main core of the paper is accompanied by appendices. Appendices D to J are only available on Zenodo (see https://zenodo.org/records/13990211).
@@ 2 Description of the input data
The DR3 data correspond to the first 34 months of the nominal mission (25 July 2014 - 28 May 2017, almost 1038 d, 2.84 y). The instrumental spectra obtained at each RVS transit are converted to proper physical spectra (reduced) by the pipeline developed by the CU6 (spectroscopic processing), see Sartoretti et al. (2022); Katz et al. (2023). Per transit, one spectrum is acquired by each of the three CCDs in a row of the RVS (Cropper et al. 2018). The RVs are measured individually on the three spectra. Three methods are used to derive the radial velocities (Sartoretti et al. 2018), each of them providing three velocities (one per CCD) and a combined RV value (over the three CCDs). The three combined RVs issued from the three methods are transmitted to the Integrator that provides a single combined value (per transit). Only this single combined value is considered here but the other intermediate RVs are used as a sanity check to validate the final combined value (see Sects. 6.4.8.2 and 6.4.8.5 of Sartoretti et al. 2022). The RVs are measured by cross-correlating the observed reduced and calibrated spectra with a theoretical/synthetic spectrum computed on the basis of various stellar atmosphere models. These so-called templates are identified through a set of three atmospheric parameters (effective temperature, surface gravity, metallicity). The selection of the template used to measure the RVs is based on these three parameters only; the alpha element content is not considered as an independent parameter. The assignation of the set of parameters to the object treated is performed on the basis of three databases. They are ordered here in the order of preference. The first possible origin is an internal compilation of ground-based catalogues. The second one is from an early run of the Gaia General Stellar Parametrizer from Spectroscopy (GSP-Spec, Recio-Blanco et al. 2023). The third one is from an early run of the Gaia General Stellar Parametrizer from Photometry (GSP-Phot, Andrae et al. 2023). Finally, if none of these sources provides the necessary parameters, the latter were estimated as previously done for DR2 (DetermineAP: Sect. 6.5 of Sartoretti et al. 2018). This process DetermineAP cross-correlates the observed spectrum with each template from a limited subset of 28 selected ones, and this for every transit. The template giving per transit the highest correlation is retained, whatever is the velocity. The final adopted template is the one that is the major contributor over all the transits. More details can be found in Sect. 3.6.1 of Katz et al. (2023). The repartition among the various origins for the data set treated here is the following: ground-based catalogues (14.4 %), GSP-Spec (33.4 %), GSP-Phot (45.6 %) and DetermineAP (6.6 %). The RV measurements from transit to transit are performed with the same template. Consequently, a template mismatch is expected to essentially introduce a systematic RV shift (for single-line objects).
@@
The measurements of the RVs were performed under the supervision of the STA Development Unit. The list of good RVs corresponding to a particular star are then analysed by the MTA to detect and separate variable RV sets from constant RV sets. For each star, the median of the RVs is computed and appears in the main catalogue. The stars with variable RVs were further considered by the global pipeline and forwarded to the CU4 (Object processing: NSS). For DR3, the adopted cutting threshold corresponds to a probability of variability of 0.99 (rv_chisq_pvalue . The composite spectra and SB2 objects were already treated separately at the level of the CU6 processing (Damerdji et al. 2024) for the determination of both RVs and are then dispatched to CU4/NSS in order to search for the corresponding orbital solutions (Damerdji et al. 2024).
@@
The non-single star (NSS) processing is intended to perform the reduction of the corresponding data from the point of view of multiple star studies. In the present paper, we consider the sole spectroscopic channel for the SB1 objects.
The input data were drawn from a list of RVs (and their 1-uncertainties) that were corrected to make reference to the barycentre of the solar system. It should be pointed out that for DR3 the rms error on the barycentric correction ( km s-1) was not propagated to the uncertainty on the individual RVs.
@@
The RVs are accompanied by a time of observations expressed in barycentric Julian days (BJD). The number of points of the time series corresponds to the number of transits for which the RV has been properly measured with success. The data could present several weaknesses such that outliers and wrong RVs due, for instance, to template mismatch errors (for the latter case, most likely the same systematic for all values within a time series). These problems and additional selection rules are further documented in Katz et al. (2023) and will not be further detailed here. It should however be noticed that some objects have been rejected at the level of the input data filtering in CU6 because they correspond to a poorly performing Astrometric Global Iterative Solution (AGIS) solution or to a large error on the along-scan field angle (see its definition in Sartoretti et al. 2018). The threshold has been chosen at mas, equivalent to about 29 km s-1. Other filtering applied to the input CU6 data is described in Sect. 4 of Katz et al. (2023). In the course of the CU6 STA/MTA validation, a bias on output RVs was noticed for stars fainter than (being 0 km s-1 at magnitude 11 and reaching 75 m s-1 at magnitude 12, Katz et al. 2023). The correction was not applied for the catalogue, nor for the transmission of the RVs to the NSS processing. The effect is quite small and should only induce a global shift of all the RVs per object along with, as a consequence, a small bias on the systemic velocities (see Sect. 3).
@@
The current astrometric solution for DR3 has been released as part of the EDR3 and contains, among others, positions for a large number of objects. The astrometric solution (AGIS) is based on a treatment assuming the objects are single, unresolved stars and thus having an image characterised by the instrumental line-spread function (for the 1D observations in the astrometric channel) or point-spread function (for the 2D observations), as described in Lindegren et al. (2021). Some objects do not behave as expected according to this assumption. The pathological behaviour could be due to anomalies, but a large contribution to this problem originates in unresolved or poorly resolved double (or multiple) objects. The separation threshold is considered to be around 04-05 (see Fabricius et al. 2021; Holl et al. 2023a). The double objects not following the behaviour for single stars are associated with a bad adjustment to the global astrometric solution and are thus accompanied by a large renormalised unit weighted error (ruwe). Lindegren et al. (2021) also associated an excess noise (impacting the position) to these objects, which is delivered in the catalogue under the name astrometric_excess_noise and which corresponds to the additional dispersion necessary to accept the concerned object in the global solution. Objects with a ruwe larger than 1.4 are certainly problematic, but smaller values already have an impact on the RVS calibration. Since the RVS is a slitless spectrometer, the calibration in wavelength of the spectra is requesting an epoch position of the object that is accurate enough. Therefore, for the objects that are deviating from the single star behaviour, we could expect that the wavelength calibration is not correct, and thus the measured transit RVs are somewhat biased and this creates artefacts in the time series. No correction of this effect is offered for DR3. As a first approach, we preferred not to discard these objects on the basis of the ruwe. In order to take into account this problem, it was decided to inflate the uncertainty on the RV value by introducing (quadratically) an additional component linked to the astrometry. This term is derived as being (km s-1) = 0.146 astrometric_excess_noise (mas) where the factor 0.146 is related to the dispersion of the spectrometer (transformation of mas into km s-1). Typically, the values of astrometric_excess_noise are below 10 mas and, thus, is below 1.46 km s-1.
@@
Some selections were also made on the objects at the moment of their entry in the NSS-SB1 processing. In the case of DR3, these kinds of stars are restricted to the range of effective temperatures 3875 K to 8125 K (mainly ((M))-K-G-F-(A) stars). No emission-line object was selected and, in principle, only stars without spectral peculiarities were permitted. The objects with a rv_renormalised_gof value greater than 4 were selected (Katz et al. 2023) to enter the NSS-SB1 processing because they were considered as clearly non-constant. The measurements of the individual (per transit) RVs are restricted to the range –1000 to +1000 km s-1. A global view of the selections applied beforehand is available in Table 18. This measurement of the individual RVs on individual spectra is only possible for bright objects. The distribution in magnitude of the treated objects is provided in Fig. 15 in the validation Sect. 9.2.1. A threshold on was selected at 12 mag. Fainter objects are considered to produce individual spectra for which the S/N ratio does not enable the derivation of a sound radial velocity. This strongly reduces the amount of selected objects. The stellar RVs are not processed by the NSS spectroscopic channel if the number of transits (data points) is less than 10 (see details in Sect. 6.3.1).
@@
When the SB2 chain (Damerdji et al. 2024) analyses an SB2 time series and is not able to derive the corresponding SB2 orbital solution, the object can be redirected to the SB1 chain. An SB1 time series is built with the RVs characterised by the smallest uncertainties. The object is then considered as an SB1 and is forwarded to the present chain. This is a second possibility of entering the SB1 channel. These objects are flagged (for more details, see Damerdji et al. 2024, and also Appendix I).